OpenStudy (winsteria):

What is the value of the expression when a = 4, b = 8, and c = 5? 3a + b ------ 2c CAN SOMEONE PLEASE HELP ME I'LL GIVE YOU A MEDAL AND FAN

OpenStudy (winsteria):

CAN SOMEONE PLEASE HELP ME

OpenStudy (itsbribro):

3(4)+8/2(5)

OpenStudy (math&ing001):

Replace a,b and c by their values on your expression and calculate.

OpenStudy (itsbribro):

PEMDAS

OpenStudy (anonymous):

plug in values \[((3\times \left( 4 \right)) + \left( 8 \right) )\div (2\times \left( 5 \right))\] That equals 2

OpenStudy (winsteria):

OKAY THANK YOU SO ITS 2 OR???

OpenStudy (math&ing001):

Yep !

OpenStudy (anonymous):

if 3 times 4 is 12, then add 8 which is 20. Then divide that by the product of 2 and 5, which is 10. 20 divided by 10 equals 2

OpenStudy (winsteria):

THANK YOU ALL SO MUCH

OpenStudy (winsteria):

CAN YOU HELP ME WITH ONE MORE OR TWO

OpenStudy (itsbribro):

sure

OpenStudy (anonymous):

Look up PEMDAS and try to do your homework on your own

OpenStudy (winsteria):

okay uh

OpenStudy (winsteria):

What is the value of the expression? 18 ÷ 6 + 9 ----------- 14 - 2 x 5

OpenStudy (winsteria):

yea so thats the question

OpenStudy (itsbribro):

what are the ? marks for?lol

OpenStudy (winsteria):

i dont know sometimes they just show up here i'll draw it

OpenStudy (itsbribro):

ok

OpenStudy (winsteria):